OpenStudy (yumyum247):

Using the compound angle formulas, determine the exact value of Cos(11 pie/6)

OpenStudy (yumyum247):

@agent0smith

OpenStudy (yumyum247):

Cos(2theta) = 2sintheta. costheta

OpenStudy (agent0smith):

I'd use one of https://i.ytimg.com/vi/SMVExXdtXAI/maxresdefault.jpg \[\large \cos ( x+y) = \cos x \cos y - \sin x \sin y\]Wait why do we need to use compound angle formulas... 11pi/6 can be done using reference angles alone.

OpenStudy (yumyum247):

because it says in the question to use the compound angle formula...

OpenStudy (agent0smith):

What an annoying question... why not give you an angle that'd make more sense to use it. I hate questions that make things pointlessly more difficult. \[\Large \frac{ 11\pi }{ 6 } = \frac{ 6 \pi }{ 6 }+ \frac{ 5 \pi }{ 6 } = \pi +\frac{ 5 \pi }{ 6 } \] \[\large \cos ( \pi+\frac{ 5 \pi }{ 6 }) = \cos \pi \cos \frac{ 5 \pi }{ 6 } - \sin \pi \sin \frac{ 5 \pi }{ 6 }\]

OpenStudy (yumyum247):

so do i have to find the related acute angle first and then use that in the formula or something?!?!?

OpenStudy (yumyum247):

btw sir i think you've used sum and difference idnentity and not double angle??!?1

OpenStudy (agent0smith):

"compound angle" doesn't specify which formula to use. It does not say to use double angle... Easiest way would be this way, I guess \[\Large \frac{ 11\pi }{ 6 } = \frac{ 12 \pi }{ 6 }- \frac{ \pi }{ 6 } = 2\pi -\frac{ \pi }{ 6 }\] then use the formula from the link earlier \[\large \cos ( 2\pi-\frac{ \pi }{ 6 }) = \cos 2\pi \cos \frac{ \pi }{ 6 } + \sin 2\pi \sin \frac{ \pi }{ 6 }=\]

OpenStudy (yumyum247):

oh sorry i thought it said double angle....g'sis sorry papa!!

OpenStudy (yumyum247):

ok thanks!!

OpenStudy (agent0smith):

I assume you can finish it from there?

OpenStudy (yumyum247):

btw how did you come up with 12 pei/6 and pie/6??

OpenStudy (agent0smith):

Because... \[\Large \frac{ 11 }{ 6} = \frac{ 12 }{ 6} - \frac{ 1 }{ 6} \]

OpenStudy (yumyum247):

no one has answered this simple question for me...before be use the compund angle formula, do we need to find the related acute angle first!?!?!?

OpenStudy (agent0smith):

No.

OpenStudy (agent0smith):

Well, sort of, you don't HAVE to, but it's easier.

OpenStudy (yumyum247):